# Actor-Critic Model Predictive Control: Differentiable Optimization meets Reinforcement Learning

### v7

@@ Abstract
An open research question in robotics is how to combine the benefits of model-free reinforcement learning (RL)—known for its strong task performance and flexibility in optimizing general reward formulations—with the robustness and online replanning capabilities of model predictive control (MPC). This paper provides an answer by introducing a new framework called Actor-Critic Model Predictive Control. The key idea is to embed a differentiable MPC within an actor-critic RL framework. This integration allows for short-term predictive optimization of control actions through MPC, while leveraging RL for end-to-end learning and exploration over longer horizons. Through various ablation studies, we expose the benefits of the proposed approach: it achieves better out-of-distribution behaviour, better robustness to changes in the dynamics and improved sample efficiency. Additionally, we conduct an empirical analysis that reveals a relationship between the critic’s learned value function and the cost function of the differentiable MPC, providing a deeper understanding of the interplay between the critic’s value and the MPC cost functions. Finally, we validate our method in a drone racing task on different tracks, in both simulation and the real world. Our method achieves the same superhuman performance as state-of-the-art model-free RL, showcasing speeds of up to 21 m/s. We show that the proposed architecture can achieve real-time control performance, learn complex behaviors via trial and error, and retain the predictive properties of the MPC to better handle out-of-distribution behavior.
@@ I Introduction
The animal brain’s exceptional ability to quickly learn and adjust to complex behaviors is one of its most remarkable traits, which remains unattained by robotic systems. This has often been attributed to the brain’s ability to make both immediate and long-term predictions about the consequences of its actions and plan accordingly [1, 2, 3]. In robotics and control theory, model-based control has demonstrated a wide array of tasks with commendable reliability [4, 5]. In particular, Model Predictive Control (MPC) has achieved notable success across various domains such as the operation of industrial chemical plants [6], control of legged robots [7], and agile flight with drones [8, 9, 10, 11]. The effectiveness of MPC stems from its innate capability for online replanning. This enables it to make decisions that optimize a system’s future states over a specified short time horizon.
@@
However, as tasks grow in complexity, model-based approaches necessitate substantial manual engineering, tailored to each specific task. This includes the careful crafting of the cost function, tuning of hyperparameters, and design of an appropriate planning strategy [10, 12]. Often, conservative assumptions about the task are made, leading to potentially sub-optimal task performance, for instance, in tasks where the dynamical system is taken to its limits [9, 10], or in applications that require discrete mode-switching [13]. Furthermore, the modular structure of model-based approaches may result in the progressive build-up of errors, accumulating in a cascading manner. This can compound inaccuracies, reinforce conservative estimations, and diminish the overall effectiveness of the system [14, 15, 16].
@@
Most recently, reinforcement-learning-based control has gained considerable traction, demonstrating human-champion level or even superhuman performance in various domains, such as board games [17], atari games [18], complex video games [19, 20], and, in the physical world, in drone racing, both state-based [14] and vision-based [21]. In particular, the results in [14] demonstrate that model-free RL can achieve superior performance compared to optimal control techniques when pushing the limits in autonomous racing. This success is mainly due to the ability of RL to directly optimize a non-differentiable objective, eliminating the need for proxy objectives in the form of a predefined reference time trajectory or continuous 3D path. More generally, RL offers a flexible approach to control by directly optimizing a feedback policy through interactions with the environment, eliminating the need for intermediate representations. Unlike trajectory optimization or standard optimal control methods, RL can handle sparse reward signals (such as minimizing crashes, energy consumption, or lap times) and can improve robustness by leveraging domain randomization across diverse simulated conditions. This combination of offline learning, real-time adaptability, and robustness against uncertainty often makes RL more scalable and effective in complex scenarios where optimal control algorithms may struggle to yield real-time optimal solutions [14].
@@
However, RL architectures are not without their own set of challenges [15, 22]. Learning a model-free, end-to-end policy without explicitly leveraging prior knowledge in the training process, such as physics or dynamic models, results in the need to learn everything from data, often resulting in millions of interactions within a simulator. While the end-to-end paradigm is attractive, it often lacks generalizability and robustness to out-of-distribution scenarios. This has resulted in hesitancy to apply end-to-end learned architectures to safety-critical applications and has fostered the development of approaches that advocate for the introduction of safety in learned pipelines [22, 23, 24].
@@
This article is an extension of [25], which introduced an architecture called Actor-Critic Model Predictive Control to bridge the gap between Reinforcement Learning and Model Predictive Control. This architecture equips the agent with a differentiable MPC [26], placed after the last layer of the actor network, as shown in Fig. 1. This differentiable MPC module provides the system with online replanning capabilities and allows the policy to predict and optimize the short-term consequences of its actions.
@@
Instead of relying on intermediate representations such as trajectories, we directly learn a map from observations to the cost function. Therefore, at deployment time, the control commands are drawn from solving an MPC, which leverages the system’s dynamics and finds the optimal solution given the current state. The differentiable MPC module, which incorporates a model of the system’s dynamics, provides the agent with prior knowledge even before any training data is received. The second component of our actor is the cost map, a deep neural network that encapsulates the dependencies between observations and the cost function of the MPC. In other words, while the differentiable MPC captures temporal variations inside its horizon, the neural cost module encodes the dependencies in relation to the observations. This architecture thereby incorporates two different time horizon scales: the MPC drives the short-term actions while the critic network manages the long-term ones. This represents a significant advantage over vanilla actor-critic RL, where the actor is typically a randomly initialized feedforward neural network with no domain-specific structure nor priors.
@@
[25] demonstrated that the AC-MPC architecture exhibits more stable behavior in out-of-distribution scenarios, highlighting its robustness and generalizability compared to its neural-network-only counterpart. Lastly, [25] showcased the applicability of the proposed approach in real-world scenarios using an agile quadrotor platform, demonstrating the feasibility and effectiveness of the AC-MPC architecture in dynamic and challenging environments.
@@
  • It improves the AC-MPC algorithm by incorporating predictions into the critic’s training process. This improvement leverages the prior knowledge of the system through the MPC predictions and incorporates them in the learning process of the value function, resulting in improved sample efficiency during training.
  • It conducts an empirical analysis that reveals a relationship between the Hessian of the learned value function and the learned cost functions of the differentiable MPC. This finding provides a deeper understanding of the interplay between the RL value and the MPC cost functions within the AC-MPC framework, offering valuable insights for future research.
  • It shows that after optimizing hyperparameters, specifically exploration, AC-MPC can leverage the prior knowledge included in the dynamics of the system and achieve better training performance than AC-MLP.
  • It shows that AC-MPC is more robust to dynamic parameter changes than the AC-MLP baseline, keeping high success rates for variations in mass, in XY inertia, and body rate limits. This improved robustness is attributed to the differentiable MPC’s access to the dynamics model.
  • It demonstrates in both simulation and real-world experiments that the proposed AC-MPC achieves superhuman performance on par with state-of-the-art model-free reinforcement learning in a drone racing task, attaining speeds of up to 21 m/s on a quadrotor platform.
Overall, the AC-MPC architecture not only advances the integration of RL and MPC but also offers practical improvements in out-of-distribution behavior, training efficiency, and real-world applicability, contributing to the development of more robust and efficient control systems.
@@ II Related Work
Model Predictive Control (MPC) in robotics. MPC has had an enormous success in robotic applications over the last decades [27, 28, 29, 30, 9, 10, 11, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44]. However, often the cost function needs to be crafted by hand by an expert, and the hyperparameters need to be tuned at deployment time [45, 46]. Additionally, in most MPC setups [10, 47, 48, 49, 50, 7, 51, 52, 9], the high-level task is first converted into a reference trajectory (planning) and then tracked by a controller (control). This decomposition of the problem into these distinct layers is greatly favored by the MPC methodology, largely due to the differentiability and continuity requirement of optimal control’s cost functions. The cost function is shaped to achieve accurate trajectory tracking and is decoupled from (and usually unrelated to) the high-level task objective. As a result, the hierarchical separation of the information between two components leads to systems that can become erratic in the presence of unmodeled dynamics. In practice, a series of conservative assumptions or approximations are required to counteract model mismatches and maintain controllability, resulting in systems that are no longer optimal [14, 15, 16].
@@
Reinforcement Learning (RL) in robotics. RL has risen as an attractive alternative to conventional controller design, achieving impressive performance that goes beyond model-based control in a variety of domains [17, 57, 19, 58, 59, 60]. RL optimizes a controller using sampled data and can handle nonlinear, nonconvex, non-differentiable and even sparse objectives, manifesting great flexibility in controller design. Compared to model-based optimal control, RL has a number of key advantages: most importantly, RL can optimize the performance objective of the task directly, removing the need for explicit intermediate representations, such as trajectories. Moreover, RL can solve complex tasks from raw, high-dimensional sensory input, without the need of a specific metric state [14]. However, RL is still far from reaching the level of robustness and generalisability of model-based control approaches when deployed in the real world, due to its brittleness when deployed in situations outside of the training distribution, and its lack of guarantees [22]. Model-free methods generally use black-box optimization methods and do not exploit the first-order gradient through the dynamics, thus cannot leverage the full advantage of the prior knowledge.
Combinations of MPC and learning. Several methods have been developed to learn cost functions for MPC [61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75], dynamics models for MPC [38, 76, 77, 78, 79, 80], or both simultaneously [81, 40, 26, 82, 83, 84, 85, 86, 87]. For example, in [62, 63], a policy search strategy is adopted that allows for learning the hyperparameters of a loss function for complex agile flight tasks. On the other hand, the works in [38, 61] use Bayesian Optimization to tune the hyperparameters and dynamics of MPC controllers for different tasks such as car racing. Similarly, in [70] the authors propose a new method called Deep Model Predictive Optimization (DMPO) that learns the update rule of an MPPI controller using RL. They evaluate their algorihtm on a real quadrotor platform and outperform state-of-the-art MPPI algorithms.
@@
In [25], we address this issue by leveraging the advantages of both differentiable MPC and model-free reinforcement learning. By equipping the actor with a differentiable MPC, our approach provides the agent with online replanning capabilities and with prior knowledge, which is a significant advantage over model-free RL, where the actor is a randomly initialized feedforward neural network. Unlike conventional MPC, our approach emphasizes robustness and adaptability, flexibly allowing for the optimization of intricate objectives through iterative exploration and refinement.
@@ III Methodology
In this section, we first present preliminaries about both Optimal Control and Reinforcement Learning, and then introduce our method.
@@ III-A Preliminaries
Consider the discrete-time dynamic system with continuous state and input spaces, $\bm{x}_{k}\in\mathcal{X}$ and $\bm{u}_{k}\in\mathcal{U}$ respectively. Let us denote the time discretized evolution of the system $f:\mathcal{X}\times\mathcal{U}\mapsto\mathcal{X}$ such that $\bm{x}_{k+1}=f(\bm{x}_{k},\bm{u}_{k})$, where the sub-index $k$ is used to denote states and inputs at time $t_{k}$. The general Optimal Control Problem considers the task of finding a control policy $\pi(\bm{x})$, a map from the current state to the optimal input, $\pi:\mathcal{X}\mapsto\mathcal{U}$, such that the cost function $J:\mathcal{X}\mapsto\mathbb{R}^{+}$ is minimized,
@@
where $k$ ranges from 00 to $N$ for $x_{k}$ and from 00 to $N-1$ for $\bm{u}_{k}$. In the following sections, we will go through different ways in which $J(\bm{x})$ can be chosen.

### v8

@@ Abstract
A key open challenge in agile quadrotor flight is how to combine the flexibility and task-level generality of model-free reinforcement learning (RL) with the structure and online replanning capabilities of model predictive control (MPC), aiming to leverage their complementary strengths in dynamic and uncertain environments. This paper provides an answer by introducing a new framework called Actor-Critic Model Predictive Control. The key idea is to embed a differentiable MPC within an actor-critic RL framework. This integration allows for short-term predictive optimization of control actions through MPC, while leveraging RL for end-to-end learning and exploration over longer horizons. Through various ablation studies, conducted in the context of agile quadrotor racing, we expose the benefits of the proposed approach: it achieves better out-of-distribution behaviour, better robustness to changes in the quadrotor’s dynamics and improved sample efficiency. Additionally, we conduct an empirical analysis using a quadrotor platform that reveals a relationship between the critic’s learned value function and the cost function of the differentiable MPC, providing a deeper understanding of the interplay between the critic’s value and the MPC cost functions. Finally, we validate our method in a drone racing task on different tracks, in both simulation and the real world. Our method achieves the same superhuman performance as state-of-the-art model-free RL, showcasing speeds of up to 21 m/s. We show that the proposed architecture can achieve real-time control performance, learn complex behaviors via trial and error, and retain the predictive properties of the MPC to better handle out-of-distribution behavior.
Supplementary Material
Video of the experiments: https://youtu.be/_qekrF4Emzg
@@ I Introduction
The animal brain’s exceptional ability to quickly learn and adjust to complex behaviors is one of its most remarkable traits, which remains unattained by robotic systems. This has often been attributed to the brain’s ability to make both immediate and long-term predictions about the consequences of its actions and plan accordingly [1, 2, 3]. In robotics and control theory, model-based control has demonstrated a wide array of tasks with commendable reliability [4, 5]. In particular, Model Predictive Control (MPC) has achieved notable success across various domains such as the operation of industrial chemical plants [6], control of legged robots [7]. In this work, we focus specifically on the domain of agile drone flight, where optimization-based approaches have recently shown a series of remarkable successes [8, 9, 10, 11]. In this context, the effectiveness of MPC stems from its innate capability for online replanning. This enables it to make decisions that optimize a system’s future states over a specified short time horizon.
@@
However, as tasks grow in complexity, model-based approaches necessitate substantial manual engineering, tailored to each specific task. This includes the careful crafting of the cost function, tuning of parameters, and design of an appropriate planning strategy [10, 12]. Often, conservative assumptions about the task are made, leading to potentially sub-optimal task performance, for instance, in tasks where the dynamical system is taken to its limits [9, 10], or in applications that require discrete mode-switching [13]. Furthermore, the modular structure of model-based approaches may result in the progressive build-up of errors, accumulating in a cascading manner. This can compound inaccuracies, reinforce conservative estimations, and diminish the overall effectiveness of the system [14, 15, 16]. Recently, a plethora of works on generative modeling approaches, particularly diffusion-based policies, have been proposed to mitigate compounding errors in sequential decision-making [17, 18, 19].
@@
These limitations have motivated a growing interest in data-driven and reinforcement-learning approaches for robotics [20, 21, 22, 23], and specifically for human-champion performance in drone racing, both state-based [14] and vision-based [24], where RL controllers can be optimized directly from task-level objectives rather than handcrafted cost terms. In particular, the results in [14] demonstrate that model-free RL can achieve superior performance compared to optimal control techniques when pushing the limits in autonomous racing. This success is mainly due to the ability of RL to directly optimize a non-differentiable objective, eliminating the need for proxy objectives in the form of a predefined reference time trajectory or continuous 3D path. More generally, RL offers a flexible approach to control by directly optimizing a feedback policy through interactions with the environment, eliminating the need for intermediate representations. Unlike trajectory optimization or standard optimal control methods, RL can handle sparse, intricate reward signals (such as minimizing crashes, energy consumption, or lap times). This combination of offline learning, combined with the use of techniques like domain randomization or curriculum learning often makes RL more scalable and effective in complex scenarios where optimal control algorithms may struggle to yield real-time optimal solutions [14].
@@
However, RL architectures are not without their own set of challenges [15, 25]. Learning a model-free, end-to-end policy without explicitly leveraging prior knowledge in the training process, such as physics or dynamic models, results in the need to learn everything from data, often resulting in millions of interactions within a simulator. Additionally, while RL approaches are generally better equipped to handle sparse rewards, they still struggle as reward sparsity increases [26, 27]. While the end-to-end paradigm is attractive, it often lacks generalizability and robustness to out-of-distribution scenarios. This has resulted in hesitancy to apply end-to-end learned architectures to safety-critical applications and has fostered the development of approaches that advocate for the introduction of safety in learned pipelines [25, 28, 29].
@@
This article is an extension of [30], which introduced an architecture called Actor-Critic Model Predictive Control to bridge the gap between Reinforcement Learning and Model Predictive Control, applied to agile quadrotor flight. This architecture equips the agent with a differentiable MPC [31], placed after the last layer of the actor network, as shown in Fig. 1. This differentiable MPC module provides the system with online replanning capabilities and allows the policy to predict and optimize the short-term consequences of its actions.
@@
Instead of relying on intermediate representations such as trajectories, we directly learn a map from observations to the cost function. Therefore, at deployment time, the control commands are drawn from solving an MPC, which leverages the quadrotor’s dynamics and finds the optimal solution given the current state. The differentiable MPC module, which incorporates a model of the quadrotor’s dynamics, provides the agent with prior knowledge even before any training data is received. The second component of our actor is the cost map, a deep neural network that encapsulates the dependencies between observations and the cost function of the MPC. In other words, while the differentiable MPC captures temporal variations inside its horizon, the neural cost module encodes the dependencies in relation to the observations. This architecture thereby incorporates two different time horizon scales: the MPC drives the short-term actions while the critic network manages the long-term ones. This represents a significant advantage over vanilla actor-critic RL, where the actor is typically a randomly initialized feedforward neural network with no domain-specific structure nor priors.
@@
[30] demonstrated that the AC-MPC architecture exhibits more stable behavior in out-of-distribution scenarios in agile quadrotor flight tasks, highlighting its generalizability compared to its neural-network-only counterpart. Lastly, [30] showcased the applicability of the proposed approach in real-world drone racing, demonstrating the feasibility and effectiveness of the AC-MPC architecture.
@@
  • Model Predictive Value Expansion (MPVE). We introduce an extension to our framework that incorporates Model Predictive Value Expansion, a critic-training scheme that re-uses the short-horizon state-action predictions produced by the differentiable MPC during every forward pass. Unlike general value expansion methods that may require a separate policy for generating rollouts, our approach efficiently improves the critic without incurring additional policy call overhead. Our analysis, performed on quadrotor systems, demonstrates that this method improves sample efficiency, with the benefits being most pronounced for small values of the TD parameter $\lambda$.
  • Relationship between value function and MPC cost. We perform a systematic analysis which reveals that the Hessian of the learned value function closely matches the quadratic cost matrices generated by the neural cost map. This empirical finding, obtained for the task of stabilizing a quadrotor system around hover, establishes a previously unknown fact, that there is a concrete relationship between what the critic learns and the optimisation landscape of the solver. This result not only provides a deeper understanding of the interplay between the RL value and the MPC cost functions within the AC-MPC framework, but also opens a path toward a deeper understanding of this type of controllers.
  • Exploration analysis We show that after optimizing hyperparameters, specifically exploration, AC-MPC can leverage the prior knowledge included in the quadrotor dynamics to achieve better training performance than AC-MLP on the drone racing task.
  • A broad set of ablations and additional comparisons: (i) we conduct an experiment where we show that, for the task of flying a quadrotor through a race track, AC-MPC is more robust to dynamic parameter changes than the AC-MLP baseline without the need of retraining the policy, keeping high success rates for variations in the quadrotor’s mass, XY inertia, and body rate limits. This improved robustness is attributed to the differentiable MPC’s access to the quadrotor’s dynamics. These results highlight AC‑MPC’s ability to cope with dynamics uncertainty, which is of paramount importance when testing in the extremely challenging real-world drone racing benchmark. (ii) we conduct an extra empirical analysis where we compare the training performance of our architecture in the drone racing task for the choice of three different matrix representations: Diagonal, Cholesky and Full Matrix. (iii) we have added an additional comparison with an adaptive state-of-the-art L1-MPC controller. This comparison with a robust controller strategy in the cases where we show the benefits of our architecture to out of distribution behaviour helps understanding the difficulty of the task and puts our results into a new perspective.
  • Extended real-world validation. We demonstrate in both simulation and real-world experiments that the proposed AC-MPC achieves superhuman performance in the extremely challenging task of drone racing, attaining speeds of up to 21 m/s on a real quadrotor platform. These results are on par with state-of-the-art model-free reinforcement learning and highlight that the strengths of AC-MPC do not come at the expense of performance.
  • Source code. We open source the AC-MPC module: https://github.com/uzh-rpg/acmpc_public
The AC-MPC architecture not only advances the integration of RL and MPC but also offers practical improvements in out-of-distribution behavior, training efficiency, and real-world applicability in the context of agile quadrotor flight, contributing to the development of more robust and efficient control systems.
@@ II Related Work
Model Predictive Control (MPC) in robotics. MPC has had an enormous success in robotic applications over the last decades [32, 33, 34, 35, 9, 10, 11, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47]. However, often the cost function needs to be crafted by hand by an expert, and the hyperparameters need to be tuned at deployment time [48, 49]. Additionally, in most MPC setups [50, 51, 52, 7, 53, 54, 9], the high-level task is first converted into a reference trajectory (planning) and then tracked by a controller (control). This decomposition of the problem into these distinct layers is greatly favored by the MPC methodology, largely due to the differentiability and continuity requirement of optimal control’s cost functions. The cost function is shaped to achieve accurate trajectory tracking and is decoupled from (and usually unrelated to) the high-level task objective. As a result, the hierarchical separation of the information between two components leads to systems that can become erratic in the presence of unmodeled dynamics. In practice, a series of conservative assumptions or approximations are required to counteract model mismatches and maintain controllability, resulting in controllers that are no longer optimal [14, 15, 16].
@@
Reinforcement Learning (RL) in robotics. RL has risen as an attractive alternative to conventional controller design, achieving impressive performance that goes beyond model-based control in a variety of domains [20, 59, 22, 60, 61, 62]. RL optimizes a controller using sampled data and revcan easily handle non-differentiable models and sparse objectives, manifesting great flexibility in controller design. Compared to model-based optimal control, RL has a number of key advantages: most importantly, RL can optimize the performance objective of the task directly, removing the need for explicit intermediate representations, such as trajectories. Moreover, RL can solve complex tasks from raw, high-dimensional sensory input, without the need of a specific metric state [14]. However, RL is still far from reaching the level of robustness and generalisability of model-based control approaches when deployed in the real world, due to its brittleness when deployed in situations outside of the training distribution, and its lack of guarantees [25]. Model-free methods generally use black-box optimization methods and do not exploit the first-order gradient through the dynamics, thus cannot leverage the full advantage of the prior knowledge.
Combinations of MPC and learning. Several methods have been developed to learn cost functions for MPC [63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76], dynamics models for MPC [42, 77, 78, 79, 80, 81, 82], or both simultaneously [83, 43, 31, 84, 85, 86, 87, 88, 89, 90, 91]. For example, in [64], a policy search strategy is adopted that allows for learning the hyperparameters of a loss function for complex agile flight tasks. On the other hand, the works in [42, 63] use Bayesian Optimization to tune the hyperparameters and dynamics of MPC controllers for different tasks such as car racing. Similarly, in [71] the authors propose a new method called Deep Model Predictive Optimization (DMPO) that learns the update rule of an MPPI controller using RL. They evaluate their algorihtm on a real quadrotor platform and outperform state-of-the-art MPPI algorithms.
@@
In [30], we address this issue by leveraging the advantages of both differentiable MPC and model-free reinforcement learning. By equipping the actor with a differentiable MPC, our approach provides the agent with online replanning capabilities and with prior knowledge, which is a significant advantage over model-free RL, where the actor is a randomly initialized feedforward neural network. Unlike conventional MPC, our approach benefits from RL training techniques like domain randomization, flexibly allowing for the optimization of intricate objectives through iterative exploration and refinement.
Real-World Drone Racing as a benchmark. Controlling a real world robot at the limits of handling is extremely hard, more so when dealing with an inherently unstable system such as a quadrotor. In order to show how challenging real-world drone racing is, in this paragraph we cronologically go through the developments in autonomous drone racing over the years. We start with polynomial trajectories being tracked by a tracking MPC controller, in the AlphaPilot [107, 8] competition. Both winning and second place used classical methods to generate polynomial minimum snap trajectories that could be tracked using a geometric or a MPC controller. However, these trajectories didn’t result in time optimal performance, since they were not driving the actuators to saturation. After this and because of this lack of optimality, the first time-optimal flight planner was developed [9]. However, the trajectories that were generated using this planner, although they resulted in the lowest time possible, couldn’t be tracked in the real hardware because the commands would be in saturation almost all of the time, and any model mistmatch, small disturbance, or state estimation error would drive the system to crash due to the lack of control authority in saturation. To allow control authority to be able to deploy the system, the authors needed to be extremely conservative with the planner, reducing the maximum thrust that the quadrotor was able to generate by up to 30%, resulting in final, real-world lap times that were not competitive with best human pilots. As a response to this, a new model-based approach, Model Predictive Contouring Control [10, 43] was developed, which was able to achieve better lap times than previous approaches. However, this approach needed to be tuned by a human in the real-world, resulting in extremely long tuning times, including tuning of parameters that govern the trade-off between path progress maximization and contouring error minimization. These parameters need to be tuned in the physical world for every different track. This approach became the state of the art of model-based drone racing, after many years of development. In a follow up publication  [14], the authors extensively compared the best MPC and RL approaches for drone racing, showing that RL approach can achieve superhuman lap times with high success rate in the real world and with significantly less engineering effort.
The history of the field demonstrates that achieving robust, superhuman performance in drone racing requires overcoming significant real-world robotic challenges. Therefore, we believe drone racing stands as a deeply challenging benchmark, perfectly suited for validating advanced control algorithms.
Our publication aims to close the gap and show that when using our architecture, an MPC controller can be synthetized that achieves the same performance, same success rates in the real world and better out-of-distribution behaviour than an RL controller in this extremely challenging benchmark.
@@ III Methodology
In this section, we first present preliminaries about model-based control and reinforcement learning, and then introduce our method.
@@ III-A Preliminaries
Consider the discrete-time dynamic syste with continuous state and input spaces, $\bm{x}_{k}\in\mathcal{X}$ and $\bm{u}_{k}\in\mathcal{U}$ respectively. Let us denote the time discretized evolution of the system $f:\mathcal{X}\times\mathcal{U}\mapsto\mathcal{X}$ such that $\bm{x}_{k+1}=f(\bm{x}_{k},\bm{u}_{k})$, where the sub-index $k$ is used to denote states and inputs at time $t_{k}$. We consider the general task of finding a control policy $\pi(\bm{x})$, which maps the current state to the optimal input, $\pi:\mathcal{X}\mapsto\mathcal{U}$, such that the cost function $J:\mathcal{X}\mapsto\mathbb{R}^{+}$ is minimized,
@@
where $k$ ranges from 0 to $N$ for $x_{k}$ and from 0 to N1N-1 for $\bm{u}_{k}$. In the following sections, we will go through different ways in which $J(\bm{x})$ can be chosen.